Circular No. 5895

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVAE 1993af, 1993ag, 1993ah M. Hamuy, Cerro Tololo Interamerican Observatory; and J. Maza,

University of Chile, report the discovery of three apparent supernovae on three 20-min unfiltered IIa-O plates taken by G. Valladares with the CTIO Curtis Schmidt telescope. SN 1993af (m_{pg} ∼ 17) was discovered by M. Wischnjewsky (University of Chile) on a plate taken Nov. 15.30 UT; the new object is located 220′′ east and 94′′ north of the nucleus of the spiral galaxy NGC 1808 (α = 5^{h}05^{m}59^{s}, δ = −37^{o}34′.7, equinox 1950.0). SN 1993ag (m_{pg} ∼ 18.5) was discovered by R. Antezana (University of Chile) on an exposure taken Nov. 15.33, and is located ∼ 5′′ west and 6′′ south of the nucleus of an anonymous galaxy at α = 10^{h}01^{m}25^{s}, δ = −35^{o}13′.1. SN 1993ah (m_{pg} ∼ 17) was discovered by Wischnjewsky on a plate taken on Nov. 16.13, the new object being located 0′′.5 west and 8′′ north of the nucleus of the lenticular galaxy ESO 471–27 (α = 23^{h}49^{m}15^{s}, δ = −28^{o}14′.6). Confirmation of all three supernovae was made by L. Ho (University of California, Berkeley) on CCD B and V images obtained on Nov. 23 with the CTIO 0.9-m telescope.

PSR 0823+26 X. Sun, Institute of High Energy Physics, Beijing; and J. Trümper,

K. Dennerl, and W. Becker, Max-Planck-Institut für Extraterrestrische Physik, report: “X-ray emission from the radio pulsar PSR 0823+26 was detected with the ROSAT PSPC at count rates of 0.0015 ± 0.0004 and 0.0009 ± 0.0003 count/s for the broad (0.8-2.0 keV) and hard (0.5-2.0 keV) bands at confidence levels of 5 and 6 σ, respectively. The separation between the x-ray and radio positions of 13′′ is compatible with the uncertainty of the PSPC position. There is no other object in the vicinity of the radio pulsar position in the SIMBAD and NED catalogues, and the Palomar Sky Survey image also shows that there is no object evident in this region. We conclude that this is the x-ray emission from the radio pulsar. The low count rate does not permit a timing analysis, but assuming a blackbody spectrum and adopting the column density (6.0 × 10^{20} cm^{−2}) and distance (380 pc) derived from the radio dispersion measure, the energy distribution of photons can be used to constrain the temperature and luminosity to values close to 0.1 keV and 1 × 10^{30} erg/s. The corresponding size of the emission area of 1 km^{2} implies that polar cap emission is dominating.”

1993 November 23 Daniel W. E. Green